OpenStudy (anonymous):

how do i solve the equation (x+6)2/3=9 ?

hartnn (hartnn):

hi @niki12 is it \(\Large (x+6)^{2/3} = 9\) ?? if so, raise both sides by the exponent of 3/2 and see what happens

OpenStudy (anonymous):

@hartnn Yes it is, would the answer be 15/2?

hartnn (hartnn):

lets see \((x+6) = \sqrt {(9)}^3 = 27\) how did you get 15/2 ??

hartnn (hartnn):

did you get how \(\sqrt{9}^3 = 27 \:\: ??\)

hartnn (hartnn):

because i am getting x =21

OpenStudy (anonymous):

The question probably is \[(x+6)*2/3 =9\] Solving that , (x+6)=27/2 x=27/2-12/2 x=15/2